OpenStudy (lexber):

(MEDAL+FAN) HELP PLEASE! Consider the functions f(x) = 2x + 1 and g(x) = x^2 − 10. What is the value of f[g(3)]? 39 -3 -7 -1

OpenStudy (knov):

it's -1.

jhonyy9 (jhonyy9):

like a first step g(3)=3^2 -10 = 9-10 = - 1 f(g(3))=f(-1)= 2*(-1) +1 = ? hope this will help you